Fill in the blanks

The frequency threshold to eject an electron from the surface of Nickel is 1.22*10^15 Hz.

The photon has a wavelength of ____ nm

This is a ___ wavelength for the photon to eject an electron.

This is a demonstration of the_____, primarily studied by Einstein which demonstrated the ___ nature of light.
Answer #1

1)

Given:

f = 1.22*10^15 s-1

use:

lambda = c/f

=(3.0*10^8 m/s)/(1.22*10^15 s-1)

= 2.459*10^-7 m

= 246 nm

Answer: 246 nm

2)

This is maximum wavelength of photon to eject an electron

Answer: maximum

3)

This is demonstration of photoelectric effect.

This explains particle nature of light

Answer: photoelectric effect, particle